OpenStudy (anonymous):

tanx sinx tanx - sinx -------- = ----------- tanx + sinx tanx - sinx prove each identity i got the answer just wanna double check it

OpenStudy (lalaly):

are u sure abt the denominator of RHS?

OpenStudy (anonymous):

my bad its: tanx sinx tanx - sinx -------- = ----------- tanx + sinx tanx sinx imma lil sleepy rii now could explain -.-

OpenStudy (anonymous):

tan^2 x sin^2 x= tan^2x (1-cos^2 x)=tan^2 x- sin^2 x now, factorise and rearrange

OpenStudy (lalaly):

lol ok on the LHS multiply top and bottom by cos x \[\frac{(tanxsinx)cosx}{(tanx+sinx)cosx}\]\[=\frac{sinx \times sinx}{sinx+sinxcosx}\]\[=\frac{\sin^2x}{sinx+sinxcosx}\]\[=\frac{1-\cos^2x}{sinx(1+cosx)}\]\[=\frac{(1-cosx) \cancel{(1+cosx)}}{sinx \cancel{(1+cosx)}}\]\[=\frac{1-cosx}{sinx}\]now multiply top and bottom by tanx\[=\frac{tanx-sinx}{sinxtanx}\]

OpenStudy (anonymous):

Thaaankksss Guyss <3 i know i sound so gay right noww.....lol

OpenStudy (lalaly):

lool ur welcome